OpenStudy (anonymous):

cos(tan^-1(sqrt 3)) Need help!!!

OpenStudy (anonymous):

what u need help with

OpenStudy (anonymous):

I don't really understand how to work the problem above.

OpenStudy (yolomcswagginsggg):

Remember that tangent is sine over cosine. This means one fraction divided by another will give √(3) When you divide a fraction by a fraction, you flip the second fraction. Since the result is √(3), you know that √(3) is in the numerator of the problem. That indicates the use of √(3)/2 as the first value. It indicates the first one because if it were the second one, it would be flipped and become 1/√(3). Since tangent is sine over cosine, think of when sine is √(3)/2: sin( π/3 ) = sin( 60° ) = √(3)/2 And cosine there is: cos( π/3 ) = cos( 60° ) = 1/2 Thus tangent is: tan( π/3 ) = tan( 60° ) = sin( 60° )/cos( 60° ) = [ √(3)/2 ] / [ 1/2 ] = [ √(3)/2 ] × [ 2/1 ] = [ √(3)/1 ] × [ 1/1 ] = √(3) ---- The exact value is π/3 radians or 60°. ---- If you need arctangent of 1/√(3) then you would know that the flipped fraction had √(3) and look for when cosine is √(3)/2.